OpenStudy (anonymous):

I have a few question I need help with please help me: squ48 over 12

OpenStudy (okra77):



OpenStudy (anonymous):

yea but how do I work it?

OpenStudy (anonymous):

48 = ? Answer this :)

OpenStudy (anonymous):

@GreenBeanDemphsey

OpenStudy (anonymous):

6*8

OpenStudy (anonymous):

k ! And 8= ?

OpenStudy (anonymous):

4*2 2*2

OpenStudy (anonymous):

K ! so 8=2*2*2means 2^3 Right ?:)

OpenStudy (anonymous):

yes, and 6=2*3

OpenStudy (anonymous):

Exacly ! And so we have : 2*3*2*2*2 Right ?:)

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

And we can say : we have : 2^4*3 Right ?:)

OpenStudy (anonymous):

yea but where did the 3 come from?

OpenStudy (anonymous):

we said 6=2*3 3 is for 6 :) k ?

OpenStudy (anonymous):

ok I got ya

OpenStudy (anonymous):

and we have : \[\sqrt{2^4*3}\] K ?

OpenStudy (anonymous):

okay 12=2*6 6=2*3

OpenStudy (anonymous):

K ! But now we should answer to : \[\sqrt{2^4*3}\] Can you answer this ?:)

OpenStudy (anonymous):

lol no I can

OpenStudy (anonymous):

cant

OpenStudy (anonymous):

We cant answer to the \[\sqrt{3}\] so only write it and for 2^4 ... can you answer to : \[\sqrt{4}\]

OpenStudy (anonymous):

@GreenBeanDemphsey